OpenStudy (anonymous):

(4c + 2)(2c - 7) = 0

OpenStudy (anonymous):

same thing as before: 

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

4c+2=0 or 2c-7=0 4c=-2 or 2c=7 c=-1/2 or c= 7/2

OpenStudy (anonymous):

@maria1234 I have been showing you these for some time. http://openstudy.com/users/maria1234#/updates/51762c94e4b0311143e444be @mujjazi please let the user answer. Dont just give out the answer, thank you.

OpenStudy (anonymous):

@some_someone Well its just the same, When you solve the whole question then you should answer it too! Sometimes they cant understand your method and find the answer.

OpenStudy (anonymous):

yeah @mujjazi i agree, sometimes... but i have been helping her with like ~6 of the same concept questions... but she still doesnt get it, but she doesnt ask questions???

OpenStudy (anonymous):

I get it! Sorry for interfering, Keep helping, Sorry again

OpenStudy (anonymous):

@some_someone you dont have to help me if you want im not trying to be mean or anything im not that smart but thanks for helping on the other ones

OpenStudy (anonymous):

@maria1234 well you should ask questions if you don't understand something :) if you don't we dont know when and on what you need help on :)

OpenStudy (anonymous):

& i do want to help, but "help" not just do work for you without you learning anything :)

OpenStudy (anonymous):

yea i know im doing my work this is the ones i need help with